# VisEscape: A Benchmark for Evaluating Exploration-driven Decision-making in Virtual Escape Rooms

### v2

@@ B.4. Prompt templates
(a) Prompts for ReAct module on initial step
@@
(b) Prompts for ReAct module and Retry
Table 10: Prompts for the ReAct module, including the initial step and subsequent steps.
@@
Table 11: Prompts for Memory module - for first run (memory construction).
@@ 3. Add any information not included in spatial memory and inspected objects that you think is necessary
Table 12: Prompts for Memory module (memory update).
@@
previous_scene_desc, previous_action, current_scene_desc
@@
<Previous Observation> : {previous_scene_desc}
@@
<Current Observation> : {current_scene_desc}
@@
Table 13: Prompts for Feedback module.
@@
Table 14: Prompts for captioning.
Appendix C. Dataset Construction Process
C.1. Asset Creation Using Trimble SketchUp
We use Trimble SketchUp to create object assets (receptacles and items) for the room escape dataset. Receptacles are designed with predefined states, such as open/closed, locked/unlocked, and pushed/pulled, enabling agent interactions. To accommodate scene dynamics, we model these predefined states and generate corresponding receptacles to reflect those states. The visualized examples of receptacles and items are shown in in Figure 14.
C.2. Scene Creation Using Autodesk Revit
  • Receptacle View: A close-up view of an receptacle accessed through the “inspect” action from the Wall View. It changes to reflect the state of receptacles.
Figure 15: Before and after applying visual rendering using Chaos Enscape. The left shows the view in Autodesk Revit before using Chaos Enscape, and the right shows the view in Chaos Enscape.
C.3. Enhancing Photo-realism Using Chaos Enscape
To enhance the realism of the dataset, we use Chaos Enscape, a rendering tool that processes scene data generated in Autodesk Revit. Chaos Enscape improves the visual fidelity of the scene, providing photorealistic object representations. This increased realism facilitates effective agent reasoning and interaction within the environment, enabling accurate simulation and analysis of escape room scenarios. Refer to Figure 15 to check the before and after application of Chaos Enscape,
@@ C.3. Enhancing Photo-realism Using Chaos Enscape
Appendix D. Human Study
Room SR(%) GC(%) Steps Duration(s)
1 75.0 92.5 41.8 231.3
2 100.0 100.0 67.3 319.5
3 50.0 87.5 71.1 319.5
4 75.0 93.8 112.0 447.0
5 100.0 100.0 77.3 274.5
6 100.0 100.0 37.0 117.0
7 100.0 100.0 46.8 225.0
8 100.0 100.0 53.0 567.3
9 75.0 90.0 52.0 567.3
10 75.0 91.7 54.0 204.5
11 50.0 79.2 37.8 231.3
12 75.0 92.9 31.3 144.0
13 100.0 100.0 41.5 199.8
14 100.0 100.0 57.3 468.3
15 100.0 100.0 36.0 151.0
16 100.0 100.0 40.8 106.0
17 25.0 96.9 63.3 488.3
18 100.0 100.0 40.5 399.3
19 75.0 91.7 43.8 378.5
20 75.0 100.0 50.8 530.3
Average 82.5 95.8 52.8 318.5
For the human study, Figures 17 and 18 illustrate the Gradio interface used to implement the game environment.
Table 15 presents the results of the human evaluation for 20 rooms, including the success rate, goal completion, number of steps, and time duration for experiments. Participants achieved a high success rates in most rooms. We recruited 20 participants for this evaluation and ensured fair compensation according to standard ethical guidelines. Each participant evaluated 4 rooms.
Figure 17: Human evaluation using the Gradio interface.
Figure 18: Human evaluation using the Gradio interface.

### v3

@@ Appendix G Human Study
(a) Prompts for Reasoning module on initial step
@@
(b) Prompts for Reasoning module and Retry
Table 14: Prompts for the Reasoning module-before action.
@@
Table 15: Prompts for memory management module - for first run (memory construction).
@@ 3. Add any information not included in spatial memory and inspected objects that you think is necessary
Table 16: Prompts for Memory management module after memory construction
@@
previous_scene, previous_action, current_scene
@@
<Previous Observation> : {previous_scene}
@@
<Current Observation> : {current_scene}
@@
Table 17: Prompts for Reasoning module-after action.
@@
Table 18: Prompts for captioning observations.
Prompt:
You are a game designer proposing ideas for an ’Escape Room Game’. I want to create an ’Escape Room Game’ in a room structure enclosed by four walls. Inside the room, there are various objects, and it’s an ’Escape Room Game’ where actions initiated from a specific object trigger other objects, allowing puzzles to be solved one by one. There are two types of objects: 1. Receptacle:
- Definition) Objects that can contain or hold other objects, such as tables, drawers, or wardrobes.
- Each Receptacle is visually distinct and has unique States. For example, a door without a lock has two States: ’open’ and ’closed’, but a door with a lock has three States: ’locked’, ’open’, and ’closed’. Also, a drawer with two compartments can have four States depending on the open/closed status of each compartment: ’opened_opened’, ’opened_closed’, ’closed_opened’, ’closed_closed’.
- Each Receptacle must be assigned to one of the four walls of the room.
- Receptacles can contain Items. However, considering realistic constraints, each Item must define the "Interactable State(s)" of the Receptacle where it can be interacted with (i.e., visible). For example, if you have a key and a wardrobe with States defined as [locked, closed, open], the key’s interactable state for that wardrobe is [open]. Interactable States can be multiple. For instance, for a drawer with two compartments, if a key exists in the second compartment, the key’s interactable states are [closed_open, open_open].
2. Item:
- Definition) Objects that can be triggers for interaction with specific Receptacles and other Items, such as keys, buttons, photos, etc. Items must be contained within a Receptacle. There are two types of Items:
- 2-a. Applicable Item: Items that can be picked up, stored in an inventory, and applied to other objects, such as keys or ID cards.
- 2-b. Quiz Item: Items that contain a puzzle, which, when correctly solved or triggered by a specific action, activates a change in another specific object within the game.
Furthermore, in addition to objects, the game requires internal logic for escaping the room. For example, a wardrobe is initially locked, but finding and applying a key from elsewhere unlocks it. A drawer is initially locked, but solving a quiz elsewhere opens it. The game logic can be very creative and diverse, but it must be defined according to the following constraints: 1. It must be limited to logic that changes the state of a Receptacle.
2. Actions related to game logic must be associated with an Item. For example, changing a drawer from a ’locked’ State to a ’closed’ State via an ’apply Key’ action is game logic. However, changing a drawer from an ’opened’ State to a ’closed’ State via a ’close’ action is not game logic.
If these two constraints are met, the in-game logic can be expressed atomically in the following format:
(receptacle_name.state_name_before_action, action_name(item_name),
receptacle_name.state_name_after_action, checkTrigger(interaction_name)) In summary, to design one escape room game, the following game components are needed:
- Approximately 5 Receptacles, all possible states defined for each object, and which wall they belong to.
- Approximately 5 Items, and which Receptacle they belong to.
- Approximately 5 game logic interactions, expressed in the atomic format described above.
Please design a game logic with following receptacles and items:
[Receptacle] - 2TierDrawer, Cupboard, Safe, Wardrobe, Desk, Door
[Receptacle] - Key1, key2, ID card, button, puzzle
Design an escape room game by creating interactions between objects and puzzles. Storytelling is not required, and the game does not need to be particularly creative or innovative.
Table 19: Prompts used for game logic design.
Variables:
image1, image2, [Optional]tried_answers
{History} (if tried_answers is given):
You have tried the following answers: {tried_answers}, but they are all wrong.
Prompt:
{image1} is a hint image for the password of the lock described in {image2}.
You should guess the password of the lock solely based on the hint image.
{History}
Respond ONLY with the your answer, no other text.
(a) Prompts for multi-image VQA without reasoning.
Variables:
image1, image2, [Optional]tried_answers
{History} (if tried_answers is given):
You have tried the following answers: {tried_answers}, but they are all wrong.
Prompt:
{image1} is a hint image for the password of the lock described in {image2}.
You should guess the password of the lock solely based on the hint image.
{History}
Before you guess, think first, and then guess the password. Your thought should be in section [Think], and your answer should be in section [Answer]. In [Answer], respond ONLY with the your answer, no other text.
[Think]
[Answer]
(b) Prompts for multi-image VQA with reasoning.
Table 20: Prompts for multi-image VQA.
Variables:
think, action
Prompt:
You are tasked with analyzing the logs of an AI agent in an escape room game.
I’ll give you a think-action pair that the AI agent has performed.
Your task is as follows:
1. Divide the agent’s reasoning chain into distinct units that fall into the categories explicitly given below:
- Observation: visually perceiving, describing, or analyzing the current scene
- Guess: making a guess or assumption based on priors or knowledge, without exact evidence acquired in the environment
- Hypothesis Formation: forming a hypothesis or idea to test in the environment, based on any evidence acquired in the environment
- Planning: Establishing or formulating action plans within the environment
- Recall: Retrieving information from history - remembering the state/location of objects, actions previously performed, and other relevant details
A unit could be:
- A single sentence
- Multiple sentences together
- Multiple units within a single sentence
2. If a unit does not fit into the above categories, create a new category called [None-<new category name>] and place it there.
3. List all units in the same order as they appear in the original text. Do not reorder or rearrange the units.
Input:
[Think]{think}
[Action]{action}
Response Format:
[Category of Unit 1]: Sentence or phrase of given [Think]
[Category of Unit 2]: Sentence or phrase of given [Think]
[Category of Unit 3]: Sentence or phrase of given [Think]
Response:
Table 21: Prompts for annotating components of reasoning.
@@ 2. Item:
Figure 23: User Interface generated by Gradio for conducting experiments for human trajectory.